# Online Language Splatting

### v1

@@ 1 Introduction
The key challenge in online 3D language mapping lies in efficiently integrating language features into 3D representations while preserving open-vocabulary capabilities. To address this, we introduce Online Language Splatting, the first framework to achieve near real-time, open-vocabulary 3D language mapping within a SLAM-GS system, eliminating the need for pre-generated language maps. Fig. 1 illustrates the proposed framework. In particular, our method addresses three core sub-challenges: (1) Real-time High-Resolution CLIP Embedding: Since offline, segment-centric CLIP feature preparation is a major runtime bottleneck, we replace it with a single-stage CLIP embedding and a Super-Resolution Decoder (SRD) module, enabling the generation of detailed, pixel-aligned CLIP maps in 18 ms per frame (Sec. 4.1). (2) Open-Vocabulary-Preserving Feature Compression in Novel Scenes: Unlike offline methods, which allow feature compression modules to be trained on the test scene, online methods rely on a pre-trained feature compressor to operate directly on unseen data. However, due to domain gaps, a single pre-trained autoencoder may struggle to maintain open-vocabulary capabilities when compressing CLIP features for online mapping. To address this online-specific generalization challenge, we introduce a two-stage autoencoder, where the second stage, an Online-Learned AutoEncoder (OLAE), dynamically adapts to the dominant data variance of the current scene. This further reduces feature dimensions while preserving critical information (Sec. 4.2). (3) Color-Language Optimization: Existing Lang-GS systems jointly optimized color and language using the same GS parameters, but these modalities inherently prefer distinct GS parameters (see Fig. 3). Prior work [35, 58, 41] jointly optimized RGB and language using shared GS parameters, which failed to achieve optimal performance for either modality. To address this, we disentangle RGB and language backpropagation paths, designing a set of separate GS parameters to effectively render high-quality outputs for both modalities (Sec. 4.3).
@@ 2.1 SLAM with Differentiable Rendering
Dense visual SLAM builds 3D maps in an online fashion, typically using classical representations such as voxel grids [4, 32, 51, 28, 27], Octrees [47, 46, 54], or point cloud [2, 15, 40, 1, 39, 52]. In recent years, differentiable rendering has gained popularity in SLAM, enabling the joint optimization of camera poses, maps, and implicit representations such as neural fields [44, 59, 38, 13, 10, 48] or explicit 3D Gaussians [26, 55, 34, 9, 14, 5, 8] with manageable computational overhead. In particular, MonoGS [26] introduces a highly efficient pipeline for online camera tracking and mapping with high-quality rendering [8], leveraging CUDA-based gradient updates to optimize camera poses. While these methods excel in rendering quality and accurate camera pose estimation, they focus solely on geometric and photometric optimization. We emphasize that our approach complements existing SLAM-GS methods [14, 34, 26] by introducing a novel capability: 3D open-vocabulary language mapping, extending the utility of SLAM-GS systems for more interactive applications.
@@ 5 Experiments
Evaluation Metrics. To assess object localization via text queries, we follow LangSplat to use mIoU and localization accuracy (Loc) on rendered language maps. Localization is considered successful if the highest-relevancy pixel falls within the ground-truth bounding box. To evaluate image rendering quality, we use PSNR, SSIM, and LPIPS, while camera pose tracking is assessed via ATE RMSE, following the evaluation protocol in [26]. Each method’s runtime is measured by per-frame processing time. Besides these known metrics, we also evaluate text-query 3D localization via Chamfer Distance (CD) and Earth Mover’s Distance (EMD) between the ground-truth and localized point sets.
@@ 5.1 Comparison with the State of the Art
Comparison to SLAM-GS SoTA Methods. We compare our method to recent SoTA SLAM-GS approaches in Table 3. As observed, despite integrating language feature mapping into a SLAM-GS system, our framework preserves novel view rendering performance and slightly improves camera localization compared to MonoGS [26], the SLAM framework on which we build. Compared to other SoTA SLAM-GS methods, our approach achieves the best overall performance in novel view rendering while incorporating language mapping. RTG-SLAM achieves superior ATE due to the inclusion of additional classical SLAM modules, which increases system complexity. Per-scene results are provided in supplemental Table 8.
@@
Runtime Analysis. Our entire network module runs at 23ms per frame on an RTX-3090 GPU, including 15ms for CLIP encoding, 2ms for super-resolution decoding, and 6ms for online compression with online training. Although the overall pipeline speed is bottlenecked by the MonoGS baseline, leading to 0.6–0.8s per frame, a much higher speed can be achieved with advancements in the SLAM-GS system. In contrast, the offline method LangSplat requires approximately 168s per frame (2.8 minutes), including 35s for SAM, 10s for post-processing, and an additional 123s per frame (amortized) for training the dense CLIP autoencoder on the testing scene. This total runtime underscores the significant computational cost of an offline approach.
@@ 5.2 Ablation Study
Super-Reso Decoder (SRD) in CLIP Embedding We analyze the impact of SRD on the Replica dataset, with results summarized in Table 1 and individual scene results provided in supplemental Table 7. We observe that SRD significantly improves both mIoU and Loc metrics from our basic online baseline. The underlying reasons for these improvements are evident through visual comparisons in Fig. 2 and Fig. 6. From Fig. 2, we can see that high-resolution language maps greatly enhance localization of small or distant objects. Fig. 6 demonstrates improvements in semantic boundaries and even helps resolve ambiguities between visually similar but different classes.
@@
As shown in Table 1, introducing OLAE even surpasses the in-domain fine-tuned single autoencoder, demonstrating its effectiveness in preserving semantic concepts upon compression. A more detailed per-scene analysis (supplemental Table 7) further reveals: Although in-domain fine-tuned pipelines tend to outperform on certain testing scenes similar to those observed during training, OLAE performs better on novel scenes and shows higher overall stability across all scenes. A more detailed ablation study on code size is provided in supplemental Sec. 12.
@@ 7 Super-Resolution Decoder Architecture
where is the predicted high-resolution feature map, and is the ground-truth feature map. The loss penalizes spatial discontinuities to ensure smoothness.
@@ 11 More Visualizations
12 Study on Code Size
@@ 13 Detailed Study on Disentanglement GS Parameters
| Disentangled Mode | Image Rendering | ||||
| Separate | Separate | Separate | PSNR | SSIM | LPIPS |
| ✗ | ✗ | ✗ | 31.23 | 0.901 | 0.197 |
| ✗ | ✓ | ✓ | 31.79 | 0.915 | 0.177 |
| ✓ | ✗ | ✗ | 31.75 | 0.918 | 0.180 |
| ✓ | ✓ | ✗ | 32.80 | 0.929 | 0.146 |
| ✓ | ✗ | ✓ | 33.57 | 0.939 | 0.118 |
| ✓ | ✓ | ✓ | 35.89 | 0.957 | 0.060 |

### v3

@@ 1 Introduction
The key challenge in online 3D language mapping lies in efficiently integrating language features into 3D representations while preserving open-vocabulary capabilities. To address this, we introduce Online Language Splatting, the first framework to achieve near real-time, open-vocabulary 3D language mapping within a SLAM-GS system, eliminating the need for pre-generated language maps. Fig. LABEL:teaser illustrates the proposed framework. In particular, our method addresses three core sub-challenges: (1) Real-time High-Resolution CLIP Embedding: Since offline, segment-centric CLIP feature preparation is a major runtime bottleneck, we replace it with a single-stage CLIP embedding and a Super-Resolution Decoder (SRD) module, enabling the generation of detailed, pixel-aligned CLIP maps in 18 ms per frame (Sec. 4.1). (2) Open-Vocabulary-Preserving Feature Compression in Novel Scenes: Unlike offline methods, which allow feature compression modules to be trained on the test scene, online methods rely on a pre-trained feature compressor to operate directly on unseen data. However, due to domain gaps, a single pre-trained autoencoder may struggle to maintain open-vocabulary capabilities when compressing CLIP features for online mapping. To address this online-specific generalization challenge, we introduce a two-stage autoencoder, where the second stage, an Online-Learned AutoEncoder (OLAE), dynamically adapts to the dominant data variance of the current scene. This further reduces feature dimensions while preserving critical information (Sec. 4.2). (3) Color-Language Optimization: Existing Lang-GS systems jointly optimized color and language using the same GS parameters, but these modalities inherently prefer distinct GS parameters (see Fig. 3). Prior work [35, 58, 41] jointly optimized RGB and language using shared GS parameters, which failed to achieve optimal performance for either modality. To address this, we disentangle RGB and language backpropagation paths, designing a set of separate GS parameters to effectively render high-quality outputs for both modalities (Sec. 4.3).
@@ 2.1 SLAM with Differentiable Rendering
Dense visual SLAM builds 3D maps in an online fashion, typically using classical representations such as voxel grids [4, 32, 51, 28, 27], Octrees [47, 46, 54], or point cloud [2, 15, 40, 1, 39, 52]. In recent years, differentiable rendering has gained popularity in SLAM, enabling the joint optimization of camera poses, maps, and implicit representations such as neural fields [44, 59, 38, 13, 10, 48] or explicit 3D Gaussians [26, 55, 34, 9, 14, 5, 8] with manageable computational overhead. In particular, MonoGS [26] introduces a highly efficient pipeline for online camera tracking and mapping with high-quality rendering [8], leveraging CUDA-based gradient updates to optimize camera poses. While these methods excel in rendering quality and accurate camera pose estimation, they focus solely on geometric and photometric optimization. We emphasize that our approach complements existing SLAM-GS methods [14, 34, 26] by introducing a novel capability: 3D open-vocabulary language mapping, extending the utility of SLAM-GS systems for more interactive applications. For further details refer to [chen2025survey3dgaussiansplatting].
@@ 5 Experiments
Evaluation Metrics. To assess object localization via text queries, we follow LangSplat to use mIoU and localization accuracy (Loc) on rendered language maps. Localization is considered successful if the highest-relevancy pixel falls within the ground-truth bounding box. To evaluate image rendering quality, we use PSNR, SSIM, and LPIPS, while camera pose tracking is assessed via ATE RMSE, following the evaluation protocol in [26]. We exclude the post-stage global refinement for strict online requirements. Each method’s runtime is measured by per-frame processing time. Besides these known metrics, we also evaluate text-query 3D localization via Chamfer Distance (CD) and Earth Mover’s Distance (EMD) between the ground-truth and localized point sets.
@@ 5.1 Comparison with the State of the Art
Comparison to SLAM-GS SoTA Methods. We compare our method to recent SoTA SLAM-GS approaches in Table 3. As observed, despite integrating language feature mapping into a SLAM-GS system, our framework preserves novel view rendering performance and localization is on par to MonoGS [26], the SLAM framework on which we build. Compared to other SoTA SLAM-GS methods, our approach achieves the best overall performance in novel view rendering while incorporating language mapping. RTG-SLAM achieves superior ATE due to the inclusion of additional classical SLAM modules, which increases system complexity. Per-scene results are provided in supplemental Table 8.
@@
Runtime Analysis. Our entire network module runs at 21ms per frame on an RTX-3090 GPU, including 15ms for CLIP encoding, 2ms for super-resolution decoding, and 6ms for online compression with online training. While the overall pipeline speed is currently bottlenecked by the MonoGS baseline—resulting in a runtime of 0.6–0.8 seconds per frame—significantly higher speeds are achievable with advancements in the SLAM-GS system. For instance, by integrating our method into Hi-SLAM [zhang2024hislam2], we achieve 7.05 FPS. In contrast, the offline method LangSplat requires approximately 168s per frame (2.8 minutes), including 35s for SAM, 10s for post-processing, and an additional 123s per frame (amortized) for training the dense CLIP autoencoder on the testing scene. This total runtime underscores the significant computational cost of an offline approach.
@@ 5.2 Ablation Study
Super-Reso Decoder (SRD) in CLIP Embedding We analyze the impact of SRD on the Replica dataset, with results summarized in Table 1 and individual scene results provided in supplemental Table 7. We observe that SRD significantly improves both mIoU and Loc metrics from our basic online baseline. The underlying reasons for these improvements are evident through visual comparisons in Fig. 2 and Fig. 6. From Fig. 2, we can see that high-resolution language maps greatly enhance localization of small or distant objects. Fig. 6 demonstrates improvements in semantic boundaries and even helps resolve ambiguities between visually similar but different classes. Cross-view consistency: In a new ablation study, we found that the generalized encoder improves cross-view consistency—without it, the average cosine similarity drops to 0.5. While frame-wise encoding alone cannot fully ensure cross-view consistency, the globally optimized 3D map from Gaussian Splatting (GS) inherently maintains consistency across views.
@@
As shown in Table 1, introducing OLAE even surpasses the in-domain fine-tuned single autoencoder, demonstrating its effectiveness in preserving semantic concepts upon compression. A more detailed per-scene analysis (supplemental Table 7) further reveals: Although in-domain fine-tuned pipelines tend to outperform on certain testing scenes similar to those observed during training, OLAE performs better on novel scenes and shows higher overall stability across all scenes. A more detailed ablation study on code size is provided in supplemental Sec. 12.1.
@@ 7 Super-Resolution Decoder Architecture
where is the predicted high-resolution feature map, and is the ground-truth feature map. The loss penalizes spatial discontinuities to ensure smoothness [7].
SRD is supervised using labels that are created from SAM-generated masks. For each image, multiple points are sampled, clustered, and refined to produce the most accurate mask, which is then propagated consistently across the labels. We train the model on the COCO and Omni datasets, leveraging hierarchical features and supervised labels to help the network learn to associate boundaries and propagate information effectively. This enables the network to produce high-quality, generalizable language features.
@@ 11 More Visualizations
| Scene | 1-stage (768 15) Offline | 2-stage (768 32 Pretrained, 32 15 Online) |
| Room0 | 0.514, 0.835 | 0.552, 0.810 |
| Room1 | 0.427, 0.839 | 0.505, 0.939 |
| Room2 | 0.396, 0.801 | 0.493, 0.824 |
| Office0 | 0.422, 0.761 | 0.433, 0.774 |
| Office1 | 0.409, 0.802 | 0.522, 0.826 |
| Mean | 0.434, 0.808 | 0.501, 0.835 |
12 Language Compression
12.1 Study on Code Size
@@
| Disentangled Mode | Image Rendering | ||||
| Separate | Separate | Separate | PSNR | SSIM | LPIPS |
| ✗ | ✗ | ✗ | 31.23 | 0.901 | 0.197 |
| ✗ | ✓ | ✓ | 31.79 | 0.915 | 0.177 |
| ✓ | ✗ | ✗ | 31.75 | 0.918 | 0.180 |
| ✓ | ✓ | ✗ | 32.80 | 0.929 | 0.146 |
| ✓ | ✗ | ✓ | 33.57 | 0.939 | 0.118 |
| ✓ | ✓ | ✓ | 35.89 | 0.957 | 0.060 |
12.2 Online Compression
Table 14 presents a comparison between single-stage and two-stage compression methods across multiple scenes. The results indicate that the two-stage method generally offers improvements in both mIOU and localization accuracy, with certain scenes like Room1 demonstrating more noticeable gains (mIOU from 0.427 to 0.505 and localization accuracy from 0.839 to 0.939). While the single-stage method is simpler and more memory-efficient, it may lose important language features due to aggressive compression (768D directly to 15D). Conversely, the two-stage method introduces an intermediate compression stage, preserving these language features but at the cost of increased complexity. Therefore, the choice between these two methods ultimately depends on the user’s specific requirements and constraints regarding accuracy versus resource utilization.